OpenStudy (anonymous):

Precal Help a. Solve this rational equation: 2x+12/x-2=x/x-2+5 . Make sure Patricia will understand the LCM step well. Also spend some time discussing why identifying extraneous solutions is important.

OpenStudy (aum):

Is this the correct equation? \[ \frac{2x+12}{x-2} = \frac{x}{x-2} + 5 \]

OpenStudy (aum):

\[ 2x + \frac{12}{x-2} = \frac{x}{x-2} + 5 = \\ \frac{2x}{1} + \frac{12}{x-2} = \frac{x}{x-2} + \frac 51 \\ \text{ } \\ \text{LCM of 1 and (x-2) in the denominator is (x-2)}. \\ \text{Multiply throughout by (x-2)}. \\ \text{ } \\ 2x(x-2) + 12 = x + 5(x-2) \\ 2x^2 - 4x + 12 = x + 5x - 10 \\ 2x^2 - 4x + 12 = 6x - 10 \\ 2x^2 - 4x + 12 - 6x + 10 = 0 \\ 2x^2 - 10x + 22 = 0 \\ x^2 - 5x + 11 = 0 \]Solve the quadratic for x.

OpenStudy (anonymous):

Thank you so much hun :)

OpenStudy (aum):

You are welcome.